# Structured state-space models are deep Wiener models

### v1

@@ 4.3 Initialization strategies
In the case of continuous-time DPLR parametrizations Gu et al. (2021) proposed to resort to the so-called HiPPO framework (Gu et al., 2020) to initialize the learnable matrices $\tilde{B}_{c}$, $\tilde{\Lambda}_{c}$, $\tilde{P}$, and $\tilde{Q}$. In particular, being it is associated to long-term memory properties, the HiPPO matrix is regarded as a suitable initialization for (11). Such an initialization is carried out by (i) building the HiPPO-LegS matrix, (ii) computing its normal and low-rank components, and (iii) applying the eigenprojection, which yields the diagonal matrix $\tilde{\Lambda}_{c}$ and the projected low-rank components $\tilde{P}$ and $\tilde{Q}$. The HiPPO framework also provides an initialization for the input matrix $\tilde{B}_{c}$. For more details, the interested reader is referred to Appendix A. The complex matrix $\tilde{C}_{c}$ and the real matrix $F$ are randomly initialized, e.g., with Xavier initialization method.
Gu et al. (2022) proposed a similar strategy to initialize the learnable parameters of the continuous-time diagonal parametrizations  (10). In particular, they suggested to initialize $\tilde{\Lambda}_{c}$ with the diagonal component of the normalized HiPPO-LegS matrix, discarding the low-rank terms $\tilde{P}$ and $\tilde{Q}$.
@@
Under (4.3), $\Gamma$ can be initialized to the identity matrix. The complex-valued matrices $\tilde{B}_{c}$ and $\tilde{C}_{c}$, and the real-valued matrix $F$ can be randomly initialized, e.g., using the Xavier method.
@@ 4.4 Comments on the parametrization strategies
For the sake of ease of exposition, the parameterizations have been introduced here in reverse chronological order. Indeed, the continuous-time DPLR parametrization of S4 has been the one that sparked a renewed research interest in SSMs (Gu et al., 2021). This strategy of parametrizing SSLs in the continuous-time domain has been proposed for two main reasons. First, it allows to resort to the HiPPO framework for the initialization of the learnable parameters. Second, it allows to simulate the learned SSMs with a sampling time potentially different from that of the training data, whereas conventional (discrete-time) RNNs would call for a new training procedure.
We point out, however, that the choice of parametrizing SSLs in the continuous-time domain implies the aliasing problem, usually overlooked in the SSMs literature. Since (7) is discretized with the sampling time $\tau$ of the available training data, one should ensure that (i) the eigenvalues are not initialized beyond the Nyquist-Shannon bandwidth $\frac{\pi}{\tau}$ and (ii) the eigenvalues remains within this bandwidth throughout the model’s training procedure. Eigenvalues beyond such bandwidth likely lead reduced modeling capabilities and poor performances in extrapolation especially when different sampling times are adopted.
In this regard, we note that the HiPPO-based initialization strategies proposed by Gu et al. (2021), Gu et al. (2022), and Smith et al. (2022) do not, in general, meet these requirements. It is indeed well-known that the frequencies of the HiPPO matrix’s eigenvalues grow quickly with $n_{\lambda}$. To mitigate this problem, one should initialize the timescale parameter $\Gamma$ so that $\tilde{A}_{c}$ has eigenvalues within the Nyquist-Shannon bandwidth. [^2] In the SSM literature $\Gamma$ is often randomly sampled from a distribution which is uniform in its log-space (Smith et al., 2022).
@@ 5 Computational efficiency of SSM
The convolutional form (5) enjoys a noteworthy computational efficiency since, as discussed in the remainder of this section, both the materialization of the filter $\mathfrak{B}_{R}$ and the convolution itself can be easily parallelized. This allows the computational cost of simulating the SSL over a $T$-long input sequence to scale as $\mathcal{O}(T\log T)$ in place of the $\mathcal{O}(T^{2})$ entailed by the iterative, sequential, application of (1) over the time axis (Smith et al., 2022), as in the case of RNNs. The algorithms allowing such a cheap computation of (15a) are now outlined.
@@ 5.1 Convolution via Parallel Scan
As proposed by Smith et al. (2022), the task of scalably computing (15a) can be addressed via Parallel Scan (Blelloch, 1990). This approach stems from the idea that intermediate blocks of $\mathfrak{B}_{R}$ can be computed in parallel and combined, allowing to scale as $\mathcal{O}(T\log_{2}T)$ times the complexity of the product $A\cdot A$. While such a product in the particular case of diagonal parametrization scales as $\mathcal{O}(n_{\lambda})$, for non-diagonal matrices its complexity is $\mathcal{O}(n_{\lambda}^{3})$, thus rapidly increasing with the state dimension. For this reason, while the parallel scan approach can, in principle, be applied to any parametrization, it has primarily been utilized for diagonal ones, see Smith et al. (2022); Orvieto et al. (2023). For more details on how parallel scan is defined and implemented to build $\mathfrak{B}_{R}$ the reader is addressed to Appendix H of Smith et al. (2022).
@@ 5.2 Convolution via FFT
$\eta_{0:T}=$ $\,\operatorname{iFFT}\Big{(}\mathcal{H}(\omega)\cdot\mathcal{U}(\omega)\Big{)},$ (16a)
$y_{t}=$ $\,\sigma({\eta}_{t})+Fu_{t},\qquad t\in\{0,...,T\}.$ (16b)
Note that both the FFT and iFFT operations scale with complexity $\mathcal{O}(T\log T)$. Gu et al. (2021) have shown that, for continuous-time DPLR-parametrized SSLs (like S4 models), the Fourier transform $\mathcal{H}(\omega)$ can be efficiently evaluated by resorting to black-box Cauchy kernels, see Appendix B.
@@ 6 Numerical example
To preliminary quantify the modeling capabilities of the described SSM architectures, we considered the well-known Silverbox benchmark (Pintelon and Schoukens, 2001, Chapter 3.4.6), for which a public dataset was made availably by Wigren and Schoukens (2013). The Silverbox is an electronic circuit that mimics the input-output behavior of a damped mechanical system with nonlinear elastic coefficient, and it has often been used for benchmarking, e.g., Wiener models (Tiels, 2015).
The training and validation data consist of ten experiments in which a multisine input with random frequency components is applied to the apparatus. Each experiment features $8192$ input-output datapoints, collected with a sampling rate of $610.35$ Hz. The validation dataset has been constructed by extracting $76$ subsequences (of length $T=512$ steps) from one of these experiments, selected randomly. From the remaining nine experiments $684$ training subsequences have been extracted.
@@ 6.1 Identification results
The training procedure was implemented in PyTorch 2.1 and is described in more detail in the accompanying code [^4] The source code will be released upon paper publication. or in Appendix C. Our hope is that the code can help speed up continued research on these architectures.
@@
S4 (Gu et al., 2021) — SSM whose layers are parametrized in the continuous-time domain (7)-(8) by a DLPR-structured state matrix (11) initialized via the HiPPO framework.
S5 (Smith et al., 2022) — SSM whose layers are parametrized by continuous-time diagonal systems (7)-(4) initialized via the HiPPO framework.
S5R — SSM whose layers are parametrized by continuous-time diagonal systems (7)-(4). Similar to S5, but initialized by random sampling of the eigenvalues, (4.3).
LRU (Orvieto et al., 2023) — SSM whose layers are parametrized by discrete-time diagonal systems (2)-(3.1).
In Table 1 the performance metrics scored by the best identified SSM are reported, and they are compared to some of those reported in the literature [^5] For a detailed accounts of the literature, the reader is referred to Andersson et al. (2019) and Maroli et al. (2019)., while in Appendix D the trajectory of the simulation error is reported. Note that, while these SSMs are more accurate than traditional Wiener models, they are still in line with those achieved by other deep learning models like TCNNs and LSTMs.
@@
At last, let us point out that training competitive SSMs requires, in general, a careful selection of the architecture hyperparameters. These not only include the number of layers and the state and output size of each layer, but also a suitable initialization strategy for the selected parametrization. The computational efficiency of these models and their reduced parameter footprint [^7] A diagonal SSL parametrized in continuous-time has $n_{\lambda}(3+2n_{y}+2n_{u})+n_{y}n_{u}$ learnable parameters, compared to the $n_{x}(4+3n_{x}+n_{u}+n_{y})+n_{y}$ learnable parameters of an LSTM layer. come at the cost of a slightly increased architecture design effort. Further testing of SSMs for nonlinear system identification are thus advisable to establish empirical design criteria.
@@ 7 Conclusions and research directions
Structured State-space Models represent an interesting approach for identifying deep Wiener models. We have in this paper summarized recent developments of SSMs within the machine learning community, attempting to disentangle their structure, parameterization, initialization, and simulation aspects. The goal was to make these models more accessible for the system identification community, to stimulate further investigation of these models for identification problems. We believe there is scope for interesting future developments and some future research directions are summarized below.
Minimal structures — Structures with fewer learnable parameters should be considered, as long as they ensure the model’s stability and allow for an efficient simulation.
Data-driven initializations — Novel strategies could exploit training data to improve and tailor the initialization of learnable parameters. For Wiener models, for example, initialization based on best linear approximators are known to boost the identified model performances.
Discrete- or continuous-time parametrizations? — Although the revived interest in SSMs originated from continuous-time parameterizations, Orvieto et al. (2023) questioned their real need, showing how similar performances can be achieved also from discrete-time parameterizations. This issue should also be further investigated in light of the aliasing problem arising in continuous-time parameterizations.
Extensive benchmarking — The performances of SSMs for nonlinear system identification should be assessed on a broader set of benchmarks, possibly featuring long-term time dependencies, for which SSMs were proposed.
@@ Appendix A HiPPO-LegS Matrix
In this appendix we provide more information on the construction of the HiPPO-LegS matrix, and how this matrix is used to initialize the $\tilde{\Lambda}_{c}$, $\tilde{P}$, and $\tilde{Q}$ matrices of the DPLR parameterization (11). The HiPPO framework is built on the idea of projecting the data into an high-dimensional latent space characterized by an orthogonal polynomial basis Gu et al. (2020). Consistently with the Koopman theory, whithin this latent state any dynamical system can be represented by an infinite-dimensional LTI system. The HiPPO framework gained popularity since both the projection operator and the latent linear representation admit a simple closed-form expression.
@@
where ${A}_{\textrm{LegS}}^{(N)}\in\mathbb{R}^{n_{x}\times n_{x}}$ is the orthonormal component, and $p\in\mathbb{R}^{n_{x}}$ some $1$-rank component. Denoting by $\big{[}{A}_{\textrm{LegS}}^{(N)}\big{]}_{a,b}$ the element of ${A}_{\textrm{LegS}}^{(N)}$ at position $(a,b)$, it holds that
@@
$V^{*}{A}_{\textrm{LegS}}V=\Lambda_{c}-\underbrace{V^{*}p}_{P}\underbrace{p^{\prime}V}_{Q^{*}}=A_{c}.$ (18)
Note that the rows of $A_{c}$ (and its components) must be re-ordered so that they fall into the structure (8). The resulting matrices can be therefore partitioned and used to initialize the learnable parameters $\tilde{\Lambda}_{c}$, $\tilde{P}$, and $\tilde{Q}$.
@@ Appendix B FFT-based convolution for DPLR parametrizations
Let $H(s)=C_{c}(sI-A_{c})^{-1}B_{c}$ be the transfer function of the continuous-time linear subsystem of (7), where the state matrix $A_{c}$ (8) is structured according to the DPLR parametrization (11). Without loss of generality we consider $\Gamma=1$. Then, applying the bilinear transformation $s=\frac{2}{\tau}\frac{1-z^{-1}}{1+z^{-1}}$ we get the transfer function of the discretized linear subsystem of (1),
$H(z)=C_{c}\left(\frac{2}{\tau}\frac{1-z^{-1}}{1+z^{-1}}I-A_{c}\right)^{-1}B_{c}.$ (20)
Computing (20) involves inverting an $n_{x}\times n_{x}$ matrix which might be impractical. Letting for compactness
$P=\begin{bmatrix}\tilde{P}\\ \operatorname{conj}(\tilde{P})\end{bmatrix},\qquad Q=\begin{bmatrix}\tilde{Q}\\ \operatorname{conj}(\tilde{Q})\end{bmatrix},$ (21)
Gu et al. (2021) noted that since $A_{c}=\Lambda_{c}-PQ^{*}$, by applying the Matrix Inversion Lemma (20) can be re-worked as
$H(z)=C_{c}\left[W(z)-W(z)P(I+Q^{*}W(z)P)^{-1}Q^{*}W(z)\right]^{-1}B_{c},$
@@
$W(z)=\left(\frac{2}{\tau}\frac{1-z^{-1}}{1+z^{-1}}I-\Lambda_{c}\right)^{-1}.$ (22b)
Note that (B) involves the inversion of an $n_{r}\times n_{r}$ matrix, where the low-rank dimension $n_{r}$ is often $1$ — as in Appendix A. Moreover, $W(z)$ can be easily computed since $\Lambda_{c}$ is diagonal. Finally, the Fourier transform is retrieved evaluating the transfer function on the unit circle,
$\mathcal{H}(\omega)=\left.H(z)\right\lvert_{z=\exp(i\omega\tau)},$ (23)
for $\omega\in\{\omega_{0},...,\omega_{T-1}\}$, where $\omega_{r}=\frac{2\pi}{\tau}\frac{r}{T}$. As shown by Gu et al. (2021), (B) and (23) can be reconducted to a black-box Cauchy kernel, and can be easily (and efficiently) implemented even for high-dimensional systems.

### v2

@@ 4.3 Initialization strategies
In the case of continuous-time DPLR parametrizations Gu et al. (2021) proposed to resort to the so-called HiPPO framework (Gu et al., 2020) to initialize the learnable matrices $\tilde{B}_{c}$, $\tilde{\Lambda}_{c}$, $\tilde{P}$, and $\tilde{Q}$. In particular, being it associated with long-term memory properties, the HiPPO matrix is regarded as a suitable initialization for (11). Such initialization is carried out by (i) building the HiPPO-LegS matrix, (ii) computing its normal and low-rank components, and (iii) applying the eigenprojection, which yields the diagonal matrix $\tilde{\Lambda}_{c}$ and the projected low-rank components $\tilde{P}$ and $\tilde{Q}$. The HiPPO framework also provides an initialization for the input matrix $\tilde{B}_{c}$. For more details, the interested reader is referred to Appendix A. The complex matrix $\tilde{C}_{c}$ and the real matrix $F$ are randomly initialized, e.g., with the Xavier method.
Gu et al. (2022) proposed a similar strategy to initialize the learnable parameters of the continuous-time diagonal parametrizations  (10). They suggested to initialize $\tilde{\Lambda}_{c}$ with the diagonal component of the normalized HiPPO-LegS matrix, discarding the low-rank terms $\tilde{P}$ and $\tilde{Q}$.
@@
Under (14), $\Gamma$ can be initialized to the identity matrix. The complex-valued matrices $\tilde{B}_{c}$ and $\tilde{C}_{c}$, and the real-valued matrix $F$ can be randomly initialized.
@@ 4.4 Comments on the parametrization strategies
For clarity purposes, the parameterizations have been introduced here in reverse chronological order. The continuous-time DPLR parametrization of S4 has indeed been the one that sparked a renewed research interest in SSMs (Gu et al., 2021). This strategy of parametrizing SSLs in the continuous-time domain has been proposed for two reasons. Firstly, it enables the HiPPO framework to be used for the weights’ initialization. Second, it allows simulation of the learned SSMs with a sampling time potentially different from that of the training data, while conventional (discrete-time) RNNs would call for a new training.
We point out, however, that the choice of parametrizing SSLs in the continuous-time domain implies the aliasing problem, usually overlooked in the SSMs literature. Since (7) is discretized with the sampling time $\tau$ of the available training data, one should ensure that (i) the eigenvalues are not initialized beyond the Nyquist-Shannon bandwidth $\frac{\pi}{\tau}$ and (ii) the eigenvalues remains within this bandwidth throughout the model’s training. Eigenvalues beyond such bandwidth likely lead to reduced modeling capabilities and poor performances in extrapolation especially when different sampling times are adopted. In this regard, we note that the HiPPO-based initialization strategies proposed by Gu et al. (2021), Gu et al. (2022), and Smith et al. (2022) do not, in general, meet these requirements. It is indeed well-known that the HiPPO matrix’s eigenvalues grow quickly with $n_{\lambda}$. To mitigate this issue, one should initialize the timescale parameter $\Gamma$ so that $\tilde{A}_{c}$ has eigenvalues within the Nyquist bandwidth. [^2] In the SSM literature $\Gamma$ is often randomly sampled from a distribution which is uniform in its log-space (Smith et al., 2022).
@@ 5 Computational efficiency of SSM
The convolutional form (16) enjoys a noteworthy computational efficiency since, as discussed in the remainder of this section, both the materialization of the filter $\mathfrak{B}_{R}$ and the convolution itself can be easily parallelized. This allows the computational cost of simulating the SSL over a $T$-long input sequence to scale as $\mathcal{O}(T\log T)$ in place of the $\mathcal{O}(T^{2})$ entailed by the iterative, sequential, application of (1) over the time axis (Smith et al., 2022). The algorithms allowing such a cheap computation of (16a) are now outlined.
@@ 5.1 Convolution via Parallel Scan
As proposed by Smith et al. (2022), the task of scalably computing (16a) can be addressed via Parallel Scan (Blelloch, 1990). This approach stems from the idea that intermediate blocks of $\mathfrak{B}_{R}$ can be computed in parallel and combined, allowing to scale as $\mathcal{O}(T\log_{2}T)$ times the complexity of the product $A\cdot A$. While such a product in the particular case of diagonal parametrization scales as $\mathcal{O}(n_{\lambda})$, for non-diagonal matrices its complexity is $\mathcal{O}(n_{\lambda}^{3})$, thus rapidly increasing with the state dimension. Thus, while the parallel scan approach can, in principle, be applied to any parametrization, it has primarily been utilized for diagonal ones, see Orvieto et al. (2023). For more details on how parallel scan is defined and implemented the reader is addressed to Smith et al. (2022), Appendix H.
Remark 3
Given the structure (2) of the SSL (1), one can instead simulate the subsystem $\tilde{x}_{k+1}=\tilde{A}\tilde{x}_{k}+\tilde{B}{u}_{k}$ and $\tilde{\eta}_{k}=\tilde{C}\tilde{x}_{k}+Du_{k}$, and then compute the output as $y_{k}=\sigma(2\,\mathfrak{Re}(\tilde{\eta}_{k}))+Fu_{k}$. This allows for further reduction of the computational burden without any approximations.
@@ 5.2 Convolution via FFT
$\tilde{\eta}_{0:T}=$ $\,\operatorname{iFFT}\big{(}\tilde{\mathcal{H}}(\omega)\cdot\mathcal{U}(\omega)\big{)},$ (17a)
${\eta_{t}}=$ ${2\,\mathfrak{Re}(\tilde{\eta}_{t}),\qquad\qquad t\in\{0,...,T\}},$ (17b)
$y_{t}=$ $\,\sigma({\eta}_{t})+Fu_{t},\qquad\,t\in\{0,...,T\}.$ (17c)
Note that both the FFT and iFFT operations scale with complexity $\mathcal{O}(T\log T)$. Gu et al. (2021) have shown that, for continuous-time DPLR-parametrized SSLs (like S4), the Fourier transform ~$\tilde{\mathcal{H}}(\omega)$ can be efficiently evaluated by resorting to black-box Cauchy kernels, see Appendix B. With this approach, one issue to pay attention to is frequency leakage due to the input’s nonperiodicity.
@@ 6 Numerical example
To preliminary quantify the modeling capabilities of the described SSM architectures, we considered the well-known Silverbox benchmark Wigren and Schoukens (2013), for which a public dataset is available. The Silverbox is an electronic circuit that mimics the input-output behavior of a damped mechanical system with a nonlinear elastic coefficient, and it has often been used for benchmarking, e.g., Wiener models (Tiels, 2015). The training and validation data consist of ten experiments in which a multisine input with random frequency components is applied to the apparatus. Each experiment features $8192$ input-output datapoints, collected with a sampling rate of $610.35$ Hz. The validation dataset has been constructed by extracting $76$ subsequences (of length $T=512$ steps) from a random experiment. From the remaining nine experiments $684$ training subsequences have been extracted.
@@ 6.1 Identification results
The training procedure was implemented in PyTorch 2.1 and is described in more detail in the accompanying code [^4] Source code: https://github.com/bonassifabio/SSM-sysid or in Appendix C. Our hope is that the code can help speed up continued research on these architectures.
@@
S4 (Gu et al., 2021) — SSM whose layers are parametrized in the continuous-time domain (7)-(8) by a DLPR-structured state matrix (11) initialized via HiPPO.
S5 (Smith et al., 2022) — SSM whose layers are parametrized by continuous-time diagonal systems (7)-(9) initialized via HiPPO.
S5R — SSM whose layers are parametrized by continuous-time diagonal systems (7)-(9). Similar to S5, but initialized by random sampling of the eigenvalues, (14).
LRU (Orvieto et al., 2023) — SSM whose layers are parametrized by discrete-time diagonal systems (2)-(4).
@@
In Table 1 the performance metrics scored by the best identified SSM are reported, and they are compared to some of those reported in the literature [^6] For detailed accounts of the literature, the reader is referred to Andersson et al. (2019) and Maroli et al. (2019)., while in Appendix D the trajectory of the simulation error is reported. Note that, although these SSMs are more accurate than traditional Wiener models, they are still in line with those achieved by other deep learning models like TCNNs and LSTMs, while featuring significantly less learnable parameters [^7] A diagonal SSL parametrized in continuous-time has $n_{\lambda}(3+2n_{y}+2n_{u})+n_{y}n_{u}$ learnable parameters, compared to the $n_{x}(4+3n_{x}+n_{u}+n_{y})+n_{y}$ learnable parameters of an LSTM layer. and high parallelizability of the training procedure.
At last, let us point out that training competitive SSMs requires, in general, a careful selection of the architecture hyperparameters. These not only include the number of layers and the state and output size of each layer, but also a suitable initialization strategy for the selected parametrization. The computational efficiency of these models and their reduced parameter footprint come at the cost of a slightly increased architecture design effort. Further testing of SSMs for nonlinear system identification is thus advisable to establish empirical design criteria.
@@ 7 Conclusions and research directions
Structured State-space Models represent an interesting approach to identifying deep Wiener models. In this paper we summarized recent developments of SSMs in the machine learning community, attempting to disentangle their structure, parameterization, initialization, and simulation aspects. The goal was to make these models more accessible for the system identification community, to stimulate further investigation of these models for identification. We believe there is scope for interesting future developments—possible research directions are summarized below.
Minimal structures  Structures with fewer learnable parameters should be considered, as long as they ensure the model’s stability and allow for an efficient simulation. In this context, it would be advisable to evaluate whether parameterizations in the transfer function domain, similar to dynoNet Forgione and Piga (2021), yield less overparametrized and more effective models.
Data-driven initializations  Novel strategies could exploit training data to improve and tailor the initialization of learnable parameters. For Wiener models, for example, initialization based on best linear approximators are known to boost the identified model performances.
Discrete- or continuous-time parametrizations?  Despite the revived interest in SSMs originated from continuous-time parameterizations, Orvieto et al. (2023) questioned their real need, showing how similar performances can be achieved also with discrete-time parameterizations. This issue should also be further investigated in light of the aliasing problem discussed in Section 4.4.
Extensive benchmarking  The performances of SSMs for nonlinear system identification should be assessed on a broader set of benchmarks, possibly featuring long-term time dependencies, for which SSMs were proposed.
@@ Appendix A HiPPO-LegS Matrix
In this appendix we provide more information on the construction of the HiPPO-LegS matrix, and how this matrix is used to initialize the $\tilde{\Lambda}_{c}$, $\tilde{P}$, and $\tilde{Q}$ matrices of the DPLR parameterization (11). The HiPPO framework is built on the idea of projecting the data into a high-dimensional latent space characterized by an orthogonal polynomial basis Gu et al. (2020). Consistently with the Koopman theory, within this latent state any dynamical system can be represented by an infinite-dimensional LTI system. The HiPPO framework gained popularity since both the projection operator and the latent linear representation admit a simple closed-form expression.
@@
where ${A}_{\textrm{LegS}}^{(N)}\in\mathbb{R}^{n_{\lambda}\times n_{\lambda}}$ is the orthonormal component, and $p\in{\mathbb{R}^{n_{\lambda}}}$ some $1$-rank component. Denoting by $\big{[}{A}_{\textrm{LegS}}^{(N)}\big{]}_{a,b}$ the element of ${A}_{\textrm{LegS}}^{(N)}$ at position $(a,b)$, it holds that
@@
$V^{*}{A}_{\textrm{LegS}}V={\tilde{\Lambda}_{c}-\underbrace{V^{*}p}_{\tilde{P}}\underbrace{p^{\prime}V}_{\tilde{Q}^{*}}:=\tilde{A}_{c}}.$ (19)
@@ Appendix B FFT-based convolution for DPLR parametrizations
Consider the transfer function $\tilde{H}(s)=\tilde{C}_{c}(sI-\tilde{A}_{c})^{-1}\tilde{B}_{c}$, where the state matrix $\tilde{A}_{c}$ (8) is structured according to the DPLR parametrization (11) and, without loss of generality, $\Gamma=1$. Gu et al. (2021) propose to compute $\tilde{\eta}_{0:T}$ as the response of $\tilde{H}(s)$, from which $\eta$ can then be computed as $\eta_{t}=2\mathfrak{Re}\big{(}\tilde{\eta}_{t}\big{)}$ for any $t\in\{0,...,T\}$.
To retrieve $\tilde{\eta}_{0:T}$ efficiently, $\tilde{H}(s)$ is first discretized via the bilinear transformation $s=\frac{2}{\tau}\frac{1-z^{-1}}{1+z^{-1}}$, which yields
$\tilde{H}(z)=\tilde{C}_{c}\left(\frac{2}{\tau}\frac{1-z^{-1}}{1+z^{-1}}I-\tilde{A}_{c}\right)^{-1}\tilde{B}_{c}.$ (21)
Computing (21) involves inverting an $n_{\lambda}\times n_{\lambda}$ matrix which might be impractical. Gu et al. (2021) noted that since $\tilde{A}_{c}=\tilde{\Lambda}_{c}-\tilde{P}\tilde{Q}^{*}$, by applying the Matrix Inversion Lemma (21) can be re-worked as
$\tilde{H}(z)=\tilde{C}_{c}\left[\tilde{W}(z)-\tilde{W}(z)\tilde{P}(I+\tilde{Q}^{*}\tilde{W}(z)\tilde{P})^{-1}\tilde{Q}^{*}\tilde{W}(z)\right]^{-1}\tilde{B}_{c},$
@@
$\tilde{W}(z)=\left(\frac{2}{\tau}\frac{1-z^{-1}}{1+z^{-1}}I-\tilde{\Lambda}_{c}\right)^{-1}.$ (22b)
Note that (22) involves the inversion of an $n_{r}\times n_{r}$ matrix, where the low-rank dimension $n_{r}$ is often $1$ — as in Appendix A. Moreover, $\tilde{W}(z)$ can be easily computed since $\tilde{\Lambda}_{c}$ is diagonal. Finally, the Fourier transform is retrieved evaluating the transfer function on the unit circle,
~${\tilde{\mathcal{H}}(\omega)}=\left.{\tilde{H}(z)}\right\lvert_{z=\exp(i\omega\tau)},$ (23)
for $\omega\in\{\omega_{0},...,\omega_{T-1}\}$, where $\omega_{r}=\frac{2\pi}{\tau}\frac{r}{T}$. As shown by Gu et al. (2021), (22) and (23) can be reconducted to a black-box Cauchy kernel, and can be easily (and efficiently) implemented even for high-dimensional systems.